# Multi-modal Stance Detection: New Datasets and Model

### v2

@@ Abstract
Stance detection is a challenging task that aims to identify public opinion from social media platforms with respect to specific targets. Previous work on stance detection largely focused on pure texts. In this paper, we study multi-modal stance detection for tweets consisting of texts and images, which are prevalent in today’s fast-growing social media platforms where people often post multi-modal messages. To this end, we create five new multi-modal stance detection datasets of different domains based on Twitter, in which each example consists of a text and an image. In addition, we propose a simple yet effective Targeted Multi-modal Prompt Tuning framework (TMPT), where target information is leveraged to learn multi-modal stance features from textual and visual modalities. Experimental results on our three benchmark datasets show that the proposed TMPT achieves state-of-the-art performance in multi-modal stance detection.
@@
Bin Liang1,2,4∗, Ang Li1,2thanks:  The first two authors contribute equally to this work., Jingqian Zhao1,2, Lin Gui5, Min Yang6, Yue Yu3, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 3 Peng Cheng Laboratory, Shenzhen, China 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 SIAT, Chinese Academy of Sciences, Shenzhen, China bin.liang@cuhk.edu.hk,  angli@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
@@ 1 Introduction
Figure 1: An example of a user expressing an “Against” stance on “Donald Trump” and a “Favor” stance on “Joe Biden” using multi-modal information.
For multi-modal stance identification, Weinzierl and Harabagiu (2023) discussed the multi-modal stance on frames of communication. However, this work ignored the consideration of stance towards a target, which is an important factor in stance detection. Therefore, aiming to push forward the research of multi-modal stance detection, we create five new datasets, in which each example consists of a specific target, a text, and an image. These datasets contain a total of 17,544 annotated data entries across 5 domains and 12 targets, including hot topics, politicians, and debates.
@@
3) A series of experiments on our datasets show that the proposed method significantly outperforms the baseline models [^1] To facilitate future research, all datasets and code are publicly available at https://github.com/Leon-Francis/Multi-Modal-Stance-Detection.
@@ 2 Related Work
Various methods based on conventional machine learning Hasan and Ng (2014); Mohammad et al. (2016) and deep learning Sun et al. (2018); Zheng et al. (2022); Li and Caragea (2023); Li et al. (2023) have been proposed to deal with the stance detection regarding a specific target. For performing stance detection in real-world scenarios, many existing methods focus on the task of zero-shot stance detectionLiu et al. (2021b); Liang et al. (2022b); Wen and Hauptmann (2023); Zhao et al. (2023).
@@ 3.1 Data Collection
We use TweetScraper [^4] https://github.com/jonbakerfish/TweetScraper to retrieve tweets with corresponding keywords of five datasets (shown in Appendix A), keeping the posts containing text in English and at least one image or video/GIF. For videos/GIFs, we retain their first frame. For posts with multiple images, we combine the text with each image to form multiple samples. Finally, we obtain 130000, 90000, 60000, 100000, and 80000 candidate examples of five datasets, respectively.
@@ 3.2 Data Annotation
To ensure consistency with previous stance detection work, we follow the guidelines of Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020) to annotate the multi-modal stance of Mtse, Mccq and Mwtwt. For Mruc and Mtwq, the annotation guidelines are shown in Appendix B. To facilitate the analysis of how visual modality affects stance detection in our constructed datasets, we also instruct annotators to indicate whether the image conveyed stance information. We invite eight experienced researchers [^5] We recruit experienced researchers who have worked on multi-modal learning over 3 years. to label the stance for each example. Each sample will be annotated by three different annotators, and the gold label is obtained by majority vote. For the disagreed results among the three annotators, we invited three additional annotators to annotate and then performed a majority vote to obtain the gold label [^6] During the annotation process, only 2.54% of the data need to be allocated to additional annotators..
@@ 3.4 Data Analysis
We finally got 17,544 well-annotated multi-modal samples across 5 domains and 12 targets. Each example consists of a text in English and an associated image. The statistics of datasets are reported in Table 1. Note that differences in label distribution between targets are common, which is also observed in other textual stance detection datasets Mohammad et al. (2016); Mutlu et al. (2020); Conforti et al. (2020). Through analyzing the indication of whether the image conveys stance information introduced in Section 3.2, and analyzing the type of each image, we obtained statistics shown in Table 2. It indicates that in our five multi-modal stance detection datasets, nearly half of the samples will incorporate stance information due to the combination of visual modality, therefore facilitating stance detection. Further, the proportion of image types varies greatly among different datasets. This complicates image comprehension for stance detection, which is also the main challenge of multi-modal stance detection.
@@ 4 Methodology
This section details our proposed targeted multi-modal prompt tuning framework (TMPT). Given a text $S$ and an image $I$, the goal of multi-modal stance detection is to identify the stance label $y$ for the specific target $t$. Therefore, to leverage the target information in this task, we design targeted multi-modal prompt tuning for both textual and visual modalities, which are utilized to prompt the pre-trained models for learning multi-modal stance features. The architecture of our TMPT is shown in Figure 2, containing four main components: 1) Textual Prompt Tuning, which encodes the input of textual modality based on the textual targeted prompt; 2) Visual Prompt Tuning, which encodes the input of visual modality based on the visual targeted prompt; 3) Multi-modal Fusion, which fuses the representations from textual and visual modalities to capture the stance features. 4) Multi-modal Stance Detection, which derives the stance label for an input example according to the multi-modal stance features.
@@ 4.1 Textual Prompt Tuning
That is, the textual targeted prompts are designed according to the targets and the purpose of stance detection. The examples of textual targeted prompts of other targets are shown in Table 3. In Section C.1, we introduced different textual prompt settings.
@@ 4.2 Visual Prompt Tuning
Following Dosovitskiy et al. (2021), we first divide the image $I$ into $r$ fixed-sized patches $I=\{{\bm{p}}_{i}\in\mathbb{R}^{l^{2}\times c}\}_{i=1}^{r}$, where $(l,l)$ is the resolution of each patch, $c$ is the number of channels. Then, following Jia et al. (2022), we embed each patch into a $d^{V}$-dimensional vector with positional encoding ${\bm{E}}_{pos}$ by the embedding projection ${\bm{E}}$:
@@ 4.4 Multi-modal Stance Detection
Textual BERT 50.97 65.29 66.57 69.56 65.03 69.74 67.99 79.21 48.44 64.32 63.44 54.32
RoBERTa 51.67 65.03 67.93 69.31 66.32 71.62 66.01 77.41 48.74 67.64 65.96 54.40
KEBERT 56.81 68.26 66.84 71.67 67.56 69.29 69.74 80.57 49.11 65.20 72.22 61.67
@@
GPT4 68.74 66.39 65.84 63.14 65.12 69.93 71.62 52.69 47.04 50.55 68.47 45.05
Visual ResNet 37.89 38.59 47.16 39.89 42.20 43.52 37.05 50.34 39.49 39.28 44.92 38.60
ViT 40.48 40.42 46.64 46.63 50.00 40.16 46.32 50.86 38.44 41.28 46.93 36.28
SwinT 39.89 40.43 48.80 46.30 46.99 41.02 47.39 51.32 39.56 41.75 48.28 39.08
Multi- modal BERT+ViT 52.88 66.29 69.32 70.30 66.17 70.11 70.02 80.81 47.93 53.71 72.25 56.81
ViLT 35.32 48.24 47.85 62.70 56.44 58.06 60.22 73.66 45.16 50.59 56.12 59.51
CLIP 53.22 65.83 63.65 70.93 67.17 67.43 70.86 79.06 52.20 64.69 69.68 59.77
@@
TMPT+CoT 74.36 76.57 75.36 75.39 69.78 74.39 73.83 81.10 63.81 71.67 77.91 66.74
Table 4: Experimental results (%) of in-target multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Second best scores are with underlines. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$.
@@ 5.1 Data Partition
For in-target multi-modal stance detection, each dataset is divided into training, development, and testing sets with a ratio of 70%:10%:20%. For zero-shot multi-modal stance detection, the training and development set is built on known target(s), and the testing set is built on unknown target(s). Since there is only one target in the Mccq dataset, we only use it for in-target multi-modal stance detection. In the Mwtwt dataset, following Conforti et al. (2020), we select four targets (CVS_AET, CI_ESRX, ANTM_CI, and AET_HUM) to perform zero-shot scenario since DIS_FOXA is not in the same domain as them.
@@ 5.3 Experimental Settings
To leverage the powerful capabilities of large language models, following Gatto et al. (2023), we also established a variant towards our TMPT: Input text, image, and chain-of-thought generated by GPT-4-Vision (TMPT+CoT). We utilize the pre-trained uncased BERT-base Devlin et al. (2019) to embed each word as a 768-dimensional embedding and employ the pre-trained ViT-base Dosovitskiy et al. (2021) to embed each image patch as a 768-dimensional embedding, i.e., $d^{T}=d^{V}=768$. The resolution of the visual patch is set to $(16,16)$. The length of visual prompt tokens is set to $c=7$. The dimensionality of hidden vectors is set to $d_{h}=768$. We use Macro F1-score to measure the model performance. The experimental results of our models are averaged over 5 runs to ensure the final reported results are statistically stable. For detailed settings of the experiments, please refer to Appendix C.
@@
Textual BERT 26.60 32.21 59.22 59.22 64.86 57.46 22.31 18.88 27.31 15.80
RoBERTa 27.23 30.57 61.67 63.93 66.28 60.93 25.10 20.53 29.32 11.46
KEBERT 26.17 31.81 59.70 62.56 63.92 55.53 24.28 23.39 28.91 20.84
@@
GPT4 70.78 68.83 57.19 60.56 65.63 69.01 45.25 49.84 65.78 53.04
Visual ResNet 25.52 29.70 23.01 24.11 25.21 25.27 23.46 27.06 27.40 25.85
ViT 28.63 29.70 24.59 28.18 34.06 33.40 19.63 21.34 29.11 24.96
SwinT 28.54 30.85 28.53 28.50 35.87 34.33 25.13 24.39 28.30 21.19
Multi-modal BERT+ViT 26.70 31.57 59.21 59.30 65.04 59.28 24.69 22.83 26.74 20.68
ViLT 28.08 29.74 38.33 46.00 55.01 48.55 21.38 23.84 23.63 19.25
CLIP 28.21 28.99 61.08 55.67 63.80 60.06 26.29 26.39 27.15 16.33
@@
TMPT+CoT 45.58 56.16 66.75 68.69 67.28 63.78 52.24 57.40 54.17 48.90
Table 5: Experimental results of zero-shot multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$.
@@ 6.1 In-target Multi-modal Stance Detection
The results of in-target multi-modal stance detection are shown in Table 4. It can be seen that our proposed TMPT outperforms existing baselines on all datasets, which verifies the effectiveness of our method. Further, TMPT+CoT performs overall better than TMPT, which demonstrates that leveraging the knowledge from large language models can improve the comprehension of textual modality, thereby achieving better performance. For uni-modal baselines, the performance of visual modality methods is unsatisfactory, while the performance of textual modality methods is much better. This indicates that the stance expression primarily resides in the textual modality. Further, models with visual modality perform better than single-text modality ones, which proves the importance of visual modality in multi-modal stance detection. This also implies how making better use of visual information is key to improving multi-modal stance detection performance. In addition, KEBERT, which integrates specific knowledge of Twitter political posts into BERTweet, achieves promising performance compared to other textual models, which indicates that exploring external target-related knowledge might improve the performance of multi-modal stance detection. Compared with BERT+ViT, ViLT, and CLIP, our TMPT achieves better performance, denoting that exploiting targeted prompt tuning can preferably leverage the target-specific multi-modal stance information, thus improving stance detection performance.

### v3

@@ Abstract
Stance detection is a challenging task that aims to identify public opinion from social media platforms with respect to specific targets. Previous work on stance detection largely focused on pure texts. In this paper, we study multi-modal stance detection for tweets consisting of texts and images, which are prevalent in today’s fast-growing social media platforms where people often post multi-modal messages. To this end, we create five new multi-modal stance detection datasets of different domains based on Twitter, in which each example consists of a text and an image. In addition, we propose a simple yet effective Targeted Multi-modal Prompt Tuning framework (TMPT), where target information is leveraged to learn multi-modal stance features from textual and visual modalities. Experimental results on our five benchmark datasets show that the proposed TMPT achieves state-of-the-art performance in multi-modal stance detection.
@@
Bin Liang1,3,4∗, Ang Li1,3thanks:  The first two authors contribute equally to this work., Jingqian Zhao1,3, Lin Gui5, Min Yang6, Yue Yu2, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China 3 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 SIAT, Chinese Academy of Sciences, Shenzhen, China bin.liang@cuhk.edu.hk,  angli@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
@@ 1 Introduction
Figure 1: An example of a user expressing an “Against” stance towards “Donald Trump” and a “Favor” stance towards “Joe Biden” using multi-modal information.
For multi-modal stance identification, Weinzierl and Harabagiu (2023) discussed the multi-modal stance towards frames of communication. Different from conventional stance detection tasks that concentrate on the stance towards several predefined targets, their focus was centered on the frames of communication within multi-modal posts. Therefore, aiming to push forward the research of multi-modal stance detection, we create five new datasets, in which each example consists of a target, a text, and an image. These datasets contain a total of 17,544 examples across 5 domains and 12 targets, including hot topics, politicians, and debates.
@@
3) A series of experiments on our datasets show that the proposed method significantly outperforms the baseline models [^1] To facilitate future research, our datasets and code are publicly available at https://github.com/Leon-Francis/Multi-Modal-Stance-Detection.
@@ 2 Related Work
Various methods based on conventional machine learning Hasan and Ng (2014); Mohammad et al. (2016) and deep learning Sun et al. (2018); Zheng et al. (2022); Li and Caragea (2023); Li et al. (2023) have been proposed to deal with the stance detection regarding a specific target. For performing stance detection in real-world scenarios, many existing methods focus on the task of zero-shot stance detection [^2] Following Allaway and McKeown (2020), the term ”zero-shot” here refers to the model’s ability to detect stance towards targets it has not encountered during training.Liu et al. (2021b); Liang et al. (2022b); Wen and Hauptmann (2023); Zhao et al. (2023).
@@ 3.1 Data Collection
We use Twitter Streaming API [^5] https://developer.twitter.com/en/products/twitter-api to collect tweets with corresponding keywords of five datasets (shown in Appendix A), keeping the posts containing text in English and at least one image or video/GIF. For videos/GIFs, we retain their first frame because the visual information contained in consecutive frames may be very similar. For posts with multiple images, we combine the text with each image to form multiple samples due to the fact that different images may contain completely different visual information. Finally, we obtain 130000, 90000, 60000, 100000, and 80000 candidate examples of five datasets, respectively.
@@ 3.2 Data Annotation
To ensure consistency with previous stance detection work, we follow the guidelines of Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020) to annotate the multi-modal stance of Mtse, Mccq and Mwtwt. For Mruc and Mtwq, the annotation guidelines are shown in Appendix B. The meaningless/noisy posts or those that do not comply with Twitter’s policies or annotation guidelines are discarded during the annotation. We invite eight experienced researchers [^6] We recruit experienced researchers who have worked on multi-modal learning over 3 years. to label the stance for each example. Each sample will be annotated by three different annotators, and the gold label is obtained by majority vote. For the disagreed results among the three annotators, we invited three additional annotators to annotate and then performed a majority vote to obtain the gold label [^7] During the annotation process, only 2.54% of the data need to be allocated to additional annotators..
@@ 3.4 Data Analysis
We finally got 17,544 well-annotated multi-modal samples across 5 domains and 12 targets. Each example consists of a text in English and an associated image. The statistics of datasets are reported in Table 1. Note that differences in label distribution between targets are common, which is also observed in other textual stance detection datasets Mohammad et al. (2016); Mutlu et al. (2020); Conforti et al. (2020). Through analyzing the indication of whether the image conveys stance information introduced in Section 3.2, and analyzing the type of each image, we obtained statistics shown in Table 2. It can be seen that nearly half of the stance information in all five datasets comes from visual modalities. Further, the proportion of image types varies greatly among different datasets. This complicates image comprehension for stance detection, which is also the main challenge of multi-modal stance detection.
@@ 4 Methodology
In this section, we introduce our proposed targeted multi-modal prompt tuning framework (TMPT) in detail. Given a text $S$ and an image $I$, the goal of multi-modal stance detection is to identify the stance label $y$ for the specific target $t$ based on $S$ and $I$. Therefore, to leverage the target information for multi-modal stance detection, we design targeted multi-modal prompt tuning for both textual and visual modalities, which are utilized to prompt the pre-trained models for learning multi-modal stance features. The architecture of our TMPT is illustrated in Figure 2, containing four main components: 1) Textual Prompt Tuning, which encodes the input of textual modality based on the textual targeted prompt; 2) Visual Prompt Tuning, which encodes the input of visual modality based on the visual targeted prompt; 3) Multi-modal Fusion, which fuses the representations from textual and visual modalities to capture the stance features. 4) Multi-modal Stance Detection, which derives the stance label for an input example according to the multi-modal stance features.
@@ 4.1 Textual Prompt Tuning
That is, the textual targeted prompts are designed according to the targets and the purpose of stance detection. The examples of textual targeted prompts are shown in Table 3. Other textual prompt settings are introduced in Section C.1.
@@ 4.2 Visual Prompt Tuning
Following Dosovitskiy et al. (2021), we first divide the image $I$ into $r$ fixed-sized patches $I=\{{\bm{p}}_{i}\in\mathbb{R}^{l^{2}\times c}\}_{i=1}^{r}$, where $(l,l)$ is the resolution of each patch, $c$ is the number of channels. Then, following Jia et al. (2022), we flatten the patches and map to $d^{V}$-dimensional vector with a trainable linear embedding projection ${\bm{E}}$. We refer to the output of this projection as the patch embeddings. Then we concatenate these patch embeddings in the sequence dimension and added the standard learnable 1D position embeddings ${\bm{E}}_{pos}$ to the patch embeddings to retain positional information:
@@ 4.4 Multi-modal Stance Detection
Task Dataset Target # Train # Valid # Test
In-target MTSE DT 1150 170 327
JB 882 128 250
[…]
CI_ESRX 628 91 180
ANTM_CI 825 114 238
AET_HUM 674 97 186
DIS_FOXA 2081 306 599
[…]
UKR 756 108 217
[…]
TOC 1349 193 386
Zero-shot MTSE DT 1114 146 1647
JB 1434 212 1260
[…]
CI_ESRX 5994 841 899
ANTM_CI 5694 804 1177
AET_HUM 5884 840 957
[…]
UKR 971 139 1081
[…]
TOC 1222 175 1928
Table 4: Statistics of the experimental data.
@@
Textual BERT 48.25 52.04 66.57 75.62 60.85 63.05 59.24 81.53 41.25 46.80 57.77 45.91
RoBERTa 58.39 60.79 66.57 69.56 65.03 69.74 67.99 79.21 39.52 57.66 55.22 48.88
KEBERT 64.50 69.81 66.84 71.67 67.56 69.29 69.74 80.57 41.55 59.01 58.15 47.75
@@
GPT4 68.74 66.39 65.84 63.14 65.12 69.93 71.62 52.69 41.64 53.76 58.05 49.81
Visual ResNet 37.89 38.59 47.16 39.89 42.20 43.52 37.05 50.34 35.10 40.00 42.02 33.94
ViT 40.48 40.42 46.64 46.63 50.00 40.16 46.32 50.86 33.31 39.87 38.63 35.53
SwinT 39.89 40.43 48.80 46.30 46.99 41.02 47.39 51.32 35.01 40.89 35.03 35.47
Multi- modal BERT+ViT 41.86 45.82 61.32 63.20 44.71 56.45 46.85 73.71 39.28 48.41 47.47 40.86
ViLT 35.32 48.24 47.85 62.70 56.44 58.06 60.22 73.66 34.62 42.41 44.43 59.51
CLIP 53.22 65.83 63.65 70.93 67.17 67.43 70.86 79.06 44.99 59.86 55.29 40.98
@@
TMPT+CoT 66.61 68.75 71.79 74.40 69.96 68.43 63.00 82.71 45.04 60.52 68.95 59.87
Table 5: Experimental results (%) of in-target multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$. The dashed line represents a separation between fine-tuned methods, non-fine-tuned LLM-type methods, and our TMPT.
@@ 5.1 Data Partition
For in-target multi-modal stance detection, each dataset is divided into training, development, and testing sets with a ratio of 7:1:2888To address the possibility of one text corresponding to multiple images, we use the Twitter ID to split datasets, thus avoiding the issue of data leakage.. For zero-shot multi-modal stance detection, the training and development set is built on known target(s), and the testing set is built on unknown target(s). Since there is only one target in the Mccq dataset, we only use it for in-target multi-modal stance detection. In the Mwtwt dataset, following Conforti et al. (2020), we select four targets (CVS_AET, CI_ESRX, ANTM_CI, and AET_HUM) to perform zero-shot scenario since DIS_FOXA is not in the same domain as them. The statistics of each dataset are shown in Table 4.
@@ 5.3 Experimental Settings
To leverage the powerful capabilities of large language models, following Gatto et al. (2023), we propose a variant of our TMPT model, named TMPT+CoT. By utilizing GPT4-Vision to generate a chain of thought from the text and image of the sample, we concatenate this chain of thought with the text to serve as the textual modality input for TMPT+CoT. The images are used as the visual modality input. This approach is employed for both the training and testing phases. We utilize the pre-trained uncased BERT-base Devlin et al. (2019) to embed each word as a 768-dimensional embedding and employ the pre-trained ViT-base Dosovitskiy et al. (2021) to embed each image patch as a 768-dimensional embedding, i.e., $d^{T}=d^{V}=768$. The resolution of the visual patch is set to $(16,16)$. The length of visual prompt tokens is set to $c=7$. The dimensionality of hidden vectors is set to $d_{h}=768$. We use Macro F1-score to measure the model performance. The experimental results of our models are averaged over 5 runs to ensure the final reported results are statistically stable. For detailed settings of the experiments, please refer to Appendix C.
@@
Textual BERT 32.52 29.97 63.55 61.30 59.18 52.89 22.01 15.45 28.04 9.57
RoBERTa 26.60 32.21 59.22 59.22 64.86 57.46 27.10 19.98 30.62 15.84
KEBERT 26.17 31.81 59.70 62.56 63.92 55.53 24.68 28.18 29.17 19.80
@@
GPT4 70.78 68.83 57.19 60.56 65.63 69.01 40.22 49.18 62.10 52.12
Visual ResNet 25.52 29.70 23.01 24.11 25.21 25.27 23.88 25.57 27.59 24.88
ViT 28.63 29.70 24.59 28.18 34.06 33.40 27.26 28.51 29.37 23.69
SwinT 28.54 30.85 28.53 28.50 35.87 34.33 25.44 24.54 27.90 19.69
Multi-modal BERT+ViT 26.70 31.57 59.21 59.30 65.04 59.28 23.33 15.21 24.76 11.70
ViLT 28.08 29.74 38.33 46.00 55.01 48.55 21.56 23.96 23.54 19.18
CLIP 28.21 28.99 61.08 55.67 63.80 60.06 25.62 27.40 27.21 15.69
@@
TMPT+CoT 54.30 58.46 67.28 63.73 64.87 54.26 48.99 51.75 45.32 43.70
Table 6: Experimental results of zero-shot multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$. The dashed line represents a separation between fine-tuned methods, non-fine-tuned LLM-type methods, and our TMPT.
@@ 6.1 In-target Multi-modal Stance Detection
The results of in-target multi-modal stance detection are shown in Table 5. It can be seen that our proposed TMPT outperforms fine-tuning based baselines on most of datasets, denoting that exploiting targeted prompt tuning can preferably leverage the target-specific multi-modal stance information, thus improving stance detection performance. Further, TMPT+CoT performs overall better than TMPT, which demonstrates that leveraging the knowledge from large language models can improve the comprehension of textual modality, thereby achieving better performance. For uni-modal baselines, the performance of visual modality methods is unsatisfactory, while the performance of textual modality methods is much better. This indicates that the stance expression primarily resides in the textual modality. Further, CLIP, which considers both textual and visual modalities performs overall better than models which only consider textual modality, which proves the importance of visual modality in multi-modal stance detection. While, in datasets like Mwtwt and Mccq, GPT4-Vision underperforms GPT-4. Our analysis of images within these datasets revealed that a noteworthy proportion of images had comparatively high complexity levels. This also implies effective use of visual information is key to improving multi-modal stance detection performance. In addition, KEBERT, which integrates specific knowledge of Twitter political posts into BERTweet, achieves promising performance compared to other textual models, which indicates that exploring external target-related knowledge might improve the performance of multi-modal stance detection.